Rising Tide, Inc. has no debt outstanding, and its financial position is given by the following data:

             

Assets (market value = book value)

$6,500,000

EBIT

$765,000

Cost of equity

8%

Stock price

$16

Shares outstanding

406,250

Tax rate

35%

                          

The firm is considering selling bonds and simultaneously repurchasing some of its stock. If it moves to a capital structure with 20% debt based on market values, its cost of equity will increase to 9% to reflect the increased risk. Bonds can be sold at a cost of 5%. Rising Tide is a no-growth firm. Hence, all its earnings are paid out as dividends. Earnings are expected to be constant over time.

a. If the company does the proposed leveraged recapitalization, what will be the new earnings per share?

b. As a creditor, you are concerned about the company’s ability to repay its debt and interest. What is the new times interest earned?
Answer #1

(a) Market Value of Assets = $ 6500000, A recapitalization of the all-equity financed firm would see the capital structure possessing 20 % of debt and the remaining 80 % of equity. Introduction of Debt in the capital structure would lead to accrual of interest tax shield benefits to the firm, thereby raising its overall value.

Interest Cost of Bonds = 5 % and Tax Rate = 35 %

Present Value of Interest Tax Shield = 0.35 x D where D is the debt raised

Therefore, Firm Value = 6500000 + 0.35D and (6500000 + 0.35D) x 0.2 = D

1300000 + 0.07D = D

1300000 = 0.93 D

D = 1300000 / 0.93 = $ 1397849.462

Therefore, Overall Firm Value = 6500000 + 1397849.462 x 0.35 = $ 6989247.312

EBIT = $ 765000

Less: Interest Expense = 1397849.462 x 0.05 = $ 69892.4731

Profit Before Tax = $ 695107.53

Less: Tax @ 35 % = 0.35 x 695107.53 = $ 243287.63

Net Income = $ 451819.9

New Price per Share Post Debt Issue = 6989247.312 / 406250 = $ 17.2043

Number of Shares Repurchased = Debt Raised / Price per Share = 1397849.462 / 17.2043 = 81250

Shares Remaining = 406250 - 81250 = 325000

New Earnings Per Share = 451819.9 / 325000 = $ 1.39

(b) New Times Interest Earned = EBIT / Interest Expense = 765000 / 69892.4731 = 10.945